eliminate theta from the equation- cosec theta-sin theta=m and cosec theta+sin theta=n

Grade:9

1 Answers

Riddhish Bhalodia
8 years ago
ok
\frac{1}{sin\theta} - sin\theta = m................................(1)

\frac{1}{sin\theta} + sin\theta = n.................................(2)

(1) + (2) gives
sin\theta = 2/(n+m)
and (2) – (1) gives
sin\theta = (n-m)/2
equating the two sines we get an equation free of theta which is
n^2 - m^2 =4